A 12N block rests on a plane inclined 15 degrees. If the block is not moving what is the force of friction acting on that block?

1
Expert's answer
2021-12-12T16:43:35-0500


Second Newton's law in projection on x-axis:


"mg\sin\alpha = F_{fr}"

where "mg = 15N" is the weight of the block, "\alpha=15\degree" is the angle and "F_{fr}" is the frictional force. Thus, obtain:


"F_{fr}=15\cdot \sin15\degree \approx 3.9N"

Answer. 3.9 N.